The length of a rectangle is 3 feet less than half the width of the rectangle. If the perimeter of the rectangle is 174 feet, find the width and the length
Answer by checkley77(12844)

L=(W/2)-3
2[(W/2)-3]+2W=174
2W/2-6+2W=174
W+2W=174+6
3W=180
W=180/3
W=60 FOR THE WIDTH.
L=60/2-3
L=30-3
L=27 FOR THE LENGTH.
PROOF:
2*27+2*60=174
54+120=174
174=174